# How Do Recommendation Models Amplify Popularity Bias? An Analysis from the Spectral Perspective

### v3

@@ A.2 Proof of Theorem 1
1) we first showcase the relations between $\mathbf{Y}$ with $\mathbf{\hat{Y}}$ under the condition in Theorem 1 and transform $cos(\bf r,\bf q_{1})$ into $cos(\bf r,\bf\dot{q}_{1})$;
2) We then derive the preliminary lower bound of the $cos(\bf r,\bf\dot{q}_{1})$ as $cos(\bf r,\bf\dot{q}_{1})\geq\frac{{{\dot{\sigma}_{1}}{{\bf{e}}^{\top}}{{\bf{\dot{p}}}_{1}}}}{{{r_{\max}}\sqrt{\zeta(2\alpha)}}}$;
3) We further utilize the property of $\mathbf{Y}$ to give the lower bound of ${{\bf{e}}^{\top}}{{\bf{\dot{p}}}_{1}}$ as ${{\bf{e}}^{\top}}{{\bf{\dot{p}}}_{1}}\geq{\dot{\sigma}_{1}}\sqrt{1-\frac{{{r_{\max}}(\zeta(\alpha)-1)}}{{\dot{\sigma}_{1}^{2}}}}$;
4) Finally, we demonstrate $\dot{\sigma}_{1}^{2}\geq{r_{\max}}$, and give $\cos({\bf{r}},{{\bf{q}}_{1}})=\cos({\bf{r}},{{\bf{\dot{q}}}_{1}})\geq\sqrt{\frac{{2-\zeta(\alpha)}}{{\zeta(2\alpha)}}}$ when $\alpha>2$.
Part 1: the spectral relation between $\mathbf{Y}$ and $\mathbf{\hat{Y}}$. Here we focus on an embedding-based model with sufficient capacity and optimize it with MSE loss  [^1] In practice, we may introduce weight decay or negative sampling for acceleration or mitigating overfitting. But they are not our focus here we simply take the original loss for theoretical analyses.: ${L_{R}}=||{\bf{Y}}-{\bf{\hat{Y}}}||_{F}^{2}$. For other losses, like BCE loss,which can be approximated to MSE Loss via Taylor expansion [35]. Based on the theorem of PCA (Principal Component Analysis) [2], the optimal ${{\bf{\hat{Y}}}}$ will have the same spectrum as the principal dimensions of ${\bf{Y}}$. That is, their principal dimensional singular value and vector match up. Because of this relation, when analyzing the principle spectral property of ${{\bf{\hat{Y}}}}$, we can instead shift our focus and look at ${\bf{Y}}$, making our job easier.
Part 2: preliminary bound of $cos(\bf r,\bf q_{1})$. Note that $\bf r$ can be written as ${\bf{r}}={{\bf{Y}}^{\top}}{\bf{e}}$, where $\bf{e}$ denotes the $n$-dimension vector filled with ones, and ${\bf{Y}}$ can be written as:
${\bf{Y}}={\bf{\dot{P}\dot{\Sigma}}}{{\bf{\dot{Q}}}^{\bf{T}}}=\sum\limits_{k=1}^{L}{{\dot{\sigma}_{k}}{{\bf{\dot{p}}}_{k}}{\bf{\dot{q}}}_{k}^{\top}}$ (10)
through SVD decomposition. We have:
$\cos({\bf{r}},{{\bf{\dot{q}}}_{\bf{1}}})=\frac{{{{\bf{e}}^{\top}}{\bf{Y}}{{\bf{\dot{q}}}_{\bf{1}}}}}{{||{\bf{r}}||}}=\frac{{{{\bf{e}}^{\top}}\sum\limits_{k=1}^{L}{{\dot{\sigma}_{k}}{{\bf{\dot{p}}}_{k}}{\bf{\dot{q}}}_{k}^{\top}}{{\bf{\dot{q}}}_{\bf{1}}}}}{{||{\bf{r}}||}}=\frac{{{\dot{\sigma}_{1}}{{\bf{e}}^{\top}}{\bf{\dot{p}}_{1}}}}{{||{\bf{r}}||}}$ (11)
@@
$||{\bf{r}}||=\sqrt{\sum\limits_{i=1}^{m}{r_{i}^{2}}}=r_{\max}\sqrt{\sum\limits_{i=1}^{m}{{i^{-2\alpha}}}}\leq r_{\max}\sqrt{\zeta(2\alpha)}$ (12)
@@
$cos(\bf r,\bf\dot{q}_{1})\geq\frac{{{\sigma_{1}}{{\bf{e}}^{\top}}{{\bf{\dot{p}}}_{1}}}}{{{r_{\max}}\sqrt{\zeta(2\alpha)}}}$ (13)
Part 3: bound of ${{\bf{e}}^{\top}}{{\bf{\dot{p}}}_{1}}$. We first demonstrate that for the matrix ${\bf{Y}}$, we can always find a non-negative principal singular vector ${{{\bf{\dot{q}}}_{{1}}}}$. Let define matrix ${\bf{Z}}={{\bf{Y}}^{\top}}{\bf{Y}}$. We can find each element in ${{z_{kl}}}$ is non-negative. Note that:
$\dot{\sigma}_{1}^{2}=\mathop{\max}\limits_{||{\bf{\dot{q}}}||=1}{{\bf{\dot{q}}}^{\top}}{{\bf{Y}}^{\top}}{\bf{Y\dot{q}}}=\mathop{\max}\limits_{||{\bf{\dot{q}}}||=1}\sum\limits_{k=1}^{m}\sum\limits_{l=1}^{m}{{\dot{q}_{k}}{z_{kl}}{\dot{q}_{l}}}$ (14)
Suppose we have a principal singular vector ${{{\bf{q^{\prime}}}}}$ with negative elements. Let positions of these negative elements be a set $S=\{{k|\mathbf{q}^{\prime}_{k}<0}\}$. We always can construct a new $m$-dimensional vector $\bf h$ whose $k$-th element be $q^{\prime}_{k}$ if $k\in S$, be $-q^{\prime}_{k}$ otherwise. we can find $\bf h$ would be non-negative and have:
$\sum\limits_{k=1}^{m}\sum\limits_{l=1}^{m}{{h_{k}}{z_{kl}}{h_{l}}}\geq\sum\limits_{k=1}^{m}\sum\limits_{l=1}^{m}{{q^{\prime}_{k}}{z_{kl}}{q^{\prime}_{l}}}=\dot{\sigma}_{1}^{2}$ (15)
Thus, we always have a non-negative principal singular vector $\bf h$.
Let $l$ be the ID of the item with the highest popularity. For convenience, here we simply assume $l=1$. Let ${{\bf{y}}_{k}}$ denote the $k$-th column of matrix $\bf Y$. Given the non-negative of ${{{\bf{\dot{q}}}_{{1}}}}$ and $y_{ui}\in\{0,1\}$, we have the relation ${{\bf{e}}^{\top}}{\bf{\dot{p}}_{1}}\geq{{\bf{y}}^{\top}_{1}}{\bf{\dot{p}}_{1}}$. According to ${\bf{Y}^{\top}}{{\bf{\dot{p}}}_{1}}=\sigma_{1}{{\bf{\dot{q}}}_{1}}$, we further have ${{\bf{y}}^{\top}_{l}}{{\bf{\dot{p}}}_{1}}=\dot{\sigma}_{1}{\dot{q}_{11}}$, where $\dot{q}_{1k}$ denotes the $k$-th element in ${{\bf{\dot{q}}}_{1}}$.
@@
$\sum\limits_{i=2}^{m}{{\bf{y}}_{j}^{\top}{{\bf{y}}_{i}}}\leq\sum\limits_{i=2}^{m}{{\bf{y}}_{i}^{\top}{{\bf{y}}_{i}}}=\sum\limits_{i=2}^{m}{{r_{i}}}\leq{r_{\max}}(\zeta(\alpha)-1)$ (16)
It means the sum of any row of the matrix $\mathbf{B}^{\top}\mathbf{B}$ is smaller than ${r_{\max}}(\zeta(\alpha)-1)$. According to the Perron-Frobenius theorem [43], we have the largest eigenvalue value of $\mathbf{B}^{\top}\mathbf{B}$ is bounded with: ${\lambda_{1}}({{\bf{B}}^{\top}}{\bf{B}})\leq{r_{\max}}(\zeta(\alpha)-1)$. Further, we have:
$\sum\limits_{i=2}^{m}{{{({\dot{q}_{1i}})}^{2}}}=\frac{1}{{\dot{\sigma}_{1}^{2}}}||{{\bf{B}}^{\top}}{{\bf{\dot{p}}}_{1}}||_{2}^{2}\leq\frac{1}{{\dot{\sigma}_{1}^{2}}}{\lambda_{\max}}({{\bf{B}}^{\top}}{\bf{B}})\leq\frac{{{r_{\max}}}}{{\dot{\sigma}_{1}^{2}}}(\zeta(\alpha)-1)$ (17)
Considering $\bf{\dot{q}}_{1}$ is the normalization term, we can derive the lower bound of the ${{\bf{e}}^{\top}}{{\bf{\dot{p}}}_{1}}$ as:
${{\bf{e}}^{\top}}{{\bf{\dot{p}}}_{1}}\geq{\dot{\sigma}_{1}}{\dot{q}_{11}}\geq{\dot{\sigma}_{1}}\sqrt{1-\frac{{{r_{\max}}}}{{\dot{\sigma}_{1}^{2}}}(\zeta(\alpha)-1)}$ (18)
Integrating Eq. (18) into Eq. (13), finally we get the lower bound of the $\cos({\bf{r}},{{\bf{\dot{q}}}_{1}})$ as:
$\cos({\bf{r}},{{\bf{\dot{q}}}_{1}})\geq\frac{{\dot{\sigma}_{1}^{2}}}{{{r_{\max}}\sqrt{\zeta(2\alpha)}}}\sqrt{1-\frac{{{r_{\max}}(\zeta(\alpha)-1)}}{{\dot{\sigma}_{1}^{2}}}}$ (19)
Part 4: demonstrating $\dot{\sigma}_{1}^{2}\geq{r_{\max}}$. Let define a matrix ${\bf{Z}}={{\bf{Y}}^{\top}}{\bf{Y}}$. Let $l$ be the ID of the item with the highest popularity. It is easily to find that the $ll$-th element in ${\bf{Z}}$ have ${z_{ll}}=r_{max}$. Let ${\bf{v}}$ be a one-hot vector whose $l$-th element is one. we have:
${r_{\max}}={{\bf{v}}^{\top}}{\bf{Zv}}\leq\mathop{\max}\limits_{||{\bf{\dot{q}}}||=1}{{\bf{\dot{q}}}^{\top}}{\bf{Z\dot{q}}}=\dot{\sigma}_{1}^{2}$ (20)
@@
$\cos({\bf{r}},{{\bf{\dot{q}}}_{1}})\geq\sqrt{\frac{{2-\zeta(\alpha)}}{{\zeta(2\alpha)}}}$ (21)
@@
$\cos({\bf{r}},{{\bf{q}}_{1}})\geq\sqrt{\frac{{2-\zeta(\alpha)}}{{\zeta(2\alpha)}}}$ (22)
A.3 Proof of Theorem 2
@@
$S=\{u\in\mathcal{U}|{\hat{y}_{ul}}>{\hat{y}_{ui}},\forall i\in\mathcal{I}/l\}$ (23)
@@
Given the alignment of $\bf r$ and $\mathbf{q}_{1}$, and the pow-law distribution of the popularity, for any item $i\in\mathcal{I}/l$, the l.h.s of Eq. (24) can be bounded by:
${\sigma_{1}}{p_{1u}}({q_{1l}}-{q_{1i}})={\sigma_{1}}{p_{1u}}\frac{{{r_{l}}-{r_{i}}}}{{||{\bf{r}}||}}\geq{\sigma_{1}}{p_{1u}}\frac{{1-{2^{-\alpha}}}}{{\sqrt{\zeta(2\alpha)}}}$ (25)
@@
$\sum\limits_{k=2}^{L}{{\sigma_{k}}{p_{ku}}({q_{ki}}-{q_{kl}})}\leq\sum\limits_{k=2}^{L}{\sqrt{2}{\sigma_{k}}}$ (26)
@@
${\sigma_{1}}{p_{1u}}\frac{{1-{2^{-\alpha}}}}{{\sqrt{\zeta(2\alpha)}}}>\sum\limits_{k=2}^{L}{\sqrt{2}{\sigma_{k}}}$ (28)
@@
$\eta$ $=\frac{1}{n}|\{u\in U|{y_{ul}}>{y_{ui}},\forall i\in\mathcal{I}/l\}|$ (29)
@@
where $\phi(a)=\sum_{u\in\mathcal{U}}{\mathbf{I}[p_{1u}>a]}$ is an inverse cumulative function calculating the number of elements $p_{1u}$ in the left principal singular vector $\bf p_{1}$ exceeding a given value $a$, and the function $\mathbf{I}[.]$ signifies an indicator function.
A.4 Theoretical Analysis on Dimension Collapse via Gradient Optimization
Here we begin by invoking the gradient dynamic theorem from [15, 48, 6] to elucidate the dimension collapse phenomenon and then develop a theorem to illustrate how the singular values impact the popularity bias in recommendations.
Theorem 3 (Trajectory of singular values (Eq. (6) in [48])).
When training an MF model via gradient flow (gradient descent with infinitesimally small learning rate), i.e., $\frac{d}{{dt}}{\bf{U}}(t)=-\frac{{dL}}{{d{\bf{U}}}}$, $\frac{d}{{dt}}{\bf{V}}(t)=-\frac{{dL}}{{d{\bf{V}}}}$, the trajectory of singular values during the learning process obeys:
$\sigma_{k}(t)=\frac{s_{k}e^{2s_{k}t}}{e^{2s_{k}t}-1+s_{k}/\sigma_{k}(0)}$ (30)

### v6

@@ A.1. Proof of Theorem 1
1) we first showcase the relations between $\mathbf{Y}$ with $\mathbf{\hat{Y}}$ under the condition in Theorem 1 and transform $cos(\mathbf{r},\mathbf{q}_{1})$ into $cos(\mathbf{r},{\dot{\mathbf{q}}}_{1})$;
2) We then derive the preliminary lower bound of the $cos(\mathbf{r},{\mathbf{q}}_{1})$ as $cos(\mathbf{r},{\mathbf{q}}_{1})\geq\frac{{{\sigma_{1}}{{\mathbf{e}}^{\top}}{{\mathbf{p}}_{1}}}}{{{r_{\max}}\sqrt{\zeta(2\alpha)}}}$;
3) We further utilize the property of $\mathbf{Y}$ to give the lower bound of ${{\mathbf{e}}^{\top}}{{\mathbf{p}}_{1}}$ as ${{\mathbf{e}}^{\top}}{{\mathbf{p}}_{1}}\geq{\sigma_{1}}\sqrt{1-\frac{{{r_{\max}}(\zeta(\alpha)-1)}}{{\sigma_{1}^{2}}}}$;
4) Finally, we demonstrate $\sigma_{1}^{2}\geq{r_{\max}}$, and give $\cos({\mathbf{r}},{{\mathbf{q}}_{1}})=\cos({\mathbf{r}},{{\mathbf{q}}_{1}})\geq\sqrt{\frac{{2-\zeta(\alpha)}}{{\zeta(2\alpha)}}}$ when $\alpha>2$.
Part 1: the spectral relation between $\mathbf{Y}$ and $\mathbf{\hat{Y}}$. Here we focus on an embedding-based model with sufficient capacity and optimize it with MSE loss  [^1] In practice, we may introduce weight decay or negative sampling for acceleration or mitigating overfitting. But they are not our focus here we simply take the original loss for theoretical analyses.: ${L_{R}}=||{\mathbf{Y}}-{\mathbf{\hat{Y}}}||_{F}^{2}$. For other losses, like BCE loss,which can be approximated to MSE Loss via Taylor expansion (Linnainmaa, 1976). Based on the theorem of PCA (Principal Component Analysis) (Abdi and Williams, 2010), the optimal ${{\mathbf{\hat{Y}}}}$ will have the same spectrum as the principal dimensions of ${\mathbf{Y}}$. That is, their principal dimensional singular value and vector match up. Specifically. ${\mathbf{Y}}$ can be written as:
(6) ${\mathbf{Y}}={\dot{\mathbf{P}}{\dot{\Sigma}}}{{\dot{\mathbf{Q}}}^{\top}=\sum\limits_{k=1}^{L}{{\dot{\sigma}_{k}}{{\dot{\mathbf{p}}}_{k}}{\dot{\mathbf{q}}}_{k}^{\top}}}$
through SVD decomposition.
Because of this relation, when analyzing the principle spectral property of ${{\mathbf{\hat{Y}}}}$, we can instead shift our focus and look at ${\mathbf{Y}}$, making our job easier.
Part 2: preliminary bound of $cos(\mathbf{r},\mathbf{q}_{1})$. Note that $\mathbf{r}$ can be written as ${\mathbf{r}}={{\mathbf{Y}}^{\top}}{\mathbf{e}}$, where $\mathbf{e}$ denotes the $n$-dimension vector filled with ones. We have:
(7) $\cos({\mathbf{r}},{{\dot{\mathbf{q}}}_{\mathbf{1}}})=\frac{{{{\mathbf{e}}^{\top}}{\mathbf{Y}}{{\dot{\mathbf{q}}}_{\mathbf{1}}}}}{{||{\mathbf{r}}||}}=\frac{{{{\mathbf{e}}^{\top}}\sum\limits_{k=1}^{L}{{\dot{\sigma}_{k}}{{\dot{\mathbf{p}}}_{k}}{\dot{\mathbf{q}}}_{k}^{\top}}{{\dot{\mathbf{q}}}_{\mathbf{1}}}}}{{||{\mathbf{r}}||}}=\frac{{{\dot{\sigma}_{1}}{{\mathbf{e}}^{\top}}{\dot{\mathbf{p}}_{1}}}}{{||{\mathbf{r}}||}}$
@@
(8) $||{\mathbf{r}}||=\sqrt{\sum\limits_{i=1}^{m}{r_{i}^{2}}}=r_{\max}\sqrt{\sum\limits_{i=1}^{m}{{i^{-2\alpha}}}}\leq r_{\max}\sqrt{\zeta(2\alpha)}$
@@
(9) $cos(\mathbf{r},{\mathbf{q}}_{1})=cos(\mathbf{r},\dot{\mathbf{q}}_{1})\geq\frac{{{\sigma_{1}}{{\mathbf{e}}^{\top}}{{\dot{\mathbf{p}}}_{1}}}}{{{r_{\max}}\sqrt{\zeta(2\alpha)}}}$
Part 3: bound of ${{\mathbf{e}}^{\top}}{{{\mathbf{p}}}_{1}}$. We first demonstrate that for the matrix ${\mathbf{Y}}$, we can always find a non-negative principal singular vector ${{{\dot{\mathbf{q}}}_{{1}}}}$. Let define matrix ${\mathbf{Z}}={{\mathbf{Y}}^{\top}}{\mathbf{Y}}$. We can find each element in ${{z_{kl}}}$ is non-negative. Note that:
(10) $\dot{\sigma}_{1}^{2}=\mathop{\max}\limits_{||{\dot{\mathbf{q}}}||=1}{{\dot{\mathbf{q}}}^{\top}}{{\mathbf{Y}}^{\top}}{\mathbf{Y}\dot{\mathbf{q}}}=\mathop{\max}\limits_{||{\dot{\mathbf{q}}}||=1}\sum\limits_{k=1}^{m}\sum\limits_{l=1}^{m}{{\dot{q}_{k}}{z_{kl}}{\dot{q}_{l}}}$
Suppose we have a principal singular vector ${{{\mathbf{q^{\prime}}}}}$ with negative elements. Let positions of these negative elements be a set $S=\{{k|\mathbf{q}^{\prime}_{k}<0}\}$. We always can construct a new $m$-dimensional vector $\mathbf{h}$ whose $k$-th element be $q^{\prime}_{k}$ if $k\in S$, be $-q^{\prime}_{k}$ otherwise. we can find $\mathbf{h}$ would be non-negative and have:
(11) $\sum\limits_{k=1}^{m}\sum\limits_{l=1}^{m}{{h_{k}}{z_{kl}}{h_{l}}}\geq\sum\limits_{k=1}^{m}\sum\limits_{l=1}^{m}{{q^{\prime}_{k}}{z_{kl}}{q^{\prime}_{l}}}=\dot{\sigma}_{1}^{2}$
Thus, we always have a non-negative principal singular vector ${{{\dot{\mathbf{q}}}_{{1}}}}$. Since $\sigma_{1}{{\dot{\mathbf{p}}}_{1}}={\mathbf{Y}^{\top}}{{\dot{\mathbf{q}}}_{1}}$, the principal singular vector ${\dot{\mathbf{p}}}_{1}$ is also non-negative.
Let $l$ be the ID of the item with the highest popularity. For convenience, here we simply assume $l=1$. Let ${{\mathbf{y}}_{k}}$ denote the $k$-th column of matrix $\mathbf{Y}$. Given the non-negative of ${{{\dot{\mathbf{p}}}_{{1}}}}$ and $y_{ui}\in\{0,1\}$, we have the relation ${{\mathbf{e}}^{\top}}{\dot{\mathbf{p}}_{1}}\geq{{\mathbf{y}}^{\top}_{1}}{\dot{\mathbf{p}}_{1}}$. According to ${\mathbf{Y}^{\top}}{{\dot{\mathbf{p}}}_{1}}=\sigma_{1}{{\dot{\mathbf{q}}}_{1}}$, we further have ${{\mathbf{y}}^{\top}_{1}}{{\dot{\mathbf{p}}}_{1}}=\dot{\sigma}_{1}{\dot{q}_{11}}$, where $\dot{q}_{1k}$ denotes the $k$-th element in ${{\dot{\mathbf{q}}}_{1}}$.
@@
(12) $\sum\limits_{i=2}^{m}{{\mathbf{y}}_{j}^{\top}{{\mathbf{y}}_{i}}}\leq\sum\limits_{i=2}^{m}{{\mathbf{y}}_{i}^{\top}{{\mathbf{y}}_{i}}}=\sum\limits_{i=2}^{m}{{r_{i}}}\leq{r_{\max}}(\zeta(\alpha)-1)$
It means the sum of any row of the matrix $\mathbf{B}^{\top}\mathbf{B}$ is smaller than ${r_{\max}}(\zeta(\alpha)-1)$. According to the Perron-Frobenius theorem (Pillai et al., 2005), we have the largest eigenvalue value of $\mathbf{B}^{\top}\mathbf{B}$ is bounded with: ${\lambda_{1}}({{\mathbf{B}}^{\top}}{\mathbf{B}})\leq{r_{\max}}(\zeta(\alpha)-1)$. Further, we have:
(13) $\sum\limits_{i=2}^{m}{{{({\dot{q}_{1i}})}^{2}}}=\frac{1}{{\dot{\sigma}_{1}^{2}}}||{{\mathbf{B}}^{\top}}{{\dot{\mathbf{p}}}_{1}}||_{2}^{2}\leq\frac{1}{{\dot{\sigma}_{1}^{2}}}{\lambda_{\max}}({{\mathbf{B}}^{\top}}{\mathbf{B}})\leq\frac{{{r_{\max}}}}{{\dot{\sigma}_{1}^{2}}}(\zeta(\alpha)-1)$
Considering $\dot{\mathbf{q}}_{1}$ is the normalization term, we can derive the lower bound of the ${{\mathbf{e}}^{\top}}{{\dot{\mathbf{p}}}_{1}}$ as:
(14) ${{\mathbf{e}}^{\top}}{{\dot{\mathbf{p}}}_{1}}\geq{\dot{\sigma}_{1}}{\dot{q}_{11}}\geq{\dot{\sigma}_{1}}\sqrt{1-\frac{{{r_{\max}}}}{{\dot{\sigma}_{1}^{2}}}(\zeta(\alpha)-1)}$
Integrating Eq. (14) into Eq. (9), finally we get the lower bound of the $\cos({\mathbf{r}},{{\dot{\mathbf{q}}}_{1}})$ as:
(15) $\cos({\mathbf{r}},{{\dot{\mathbf{q}}}_{1}})\geq\frac{{\dot{\sigma}_{1}^{2}}}{{{r_{\max}}\sqrt{\zeta(2\alpha)}}}\sqrt{1-\frac{{{r_{\max}}(\zeta(\alpha)-1)}}{{\dot{\sigma}_{1}^{2}}}}$
Part 4: demonstrating $\dot{\sigma}_{1}^{2}\geq{r_{\max}}$. Let define a matrix ${\mathbf{Z}}={{\mathbf{Y}}^{\top}}{\mathbf{Y}}$. Let $l$ be the ID of the item with the highest popularity. It is easily to find that the $ll$-th element in ${\mathbf{Z}}$ have ${z_{ll}}=r_{max}$. Let ${\mathbf{v}}$ be a one-hot vector whose $l$-th element is one. we have:
(16) ${r_{\max}}={{\mathbf{v}}^{\top}}{\mathbf{Zv}}\leq\mathop{\max}\limits_{||{\dot{\mathbf{q}}}||=1}{{\dot{\mathbf{q}}}^{\top}}{\mathbf{Z}\dot{\mathbf{q}}}=\dot{\sigma}_{1}^{2}$
@@
(17) $\cos({\mathbf{r}},{{\dot{\mathbf{q}}}_{1}})\geq\sqrt{\frac{{2-\zeta(\alpha)}}{{\zeta(2\alpha)}}}$
@@
(18) $\cos({\mathbf{r}},{{\mathbf{q}}_{1}})\geq\sqrt{\frac{{2-\zeta(\alpha)}}{{\zeta(2\alpha)}}}$
A.2. Proof of Theorem 2
@@
(19) $S=\{u\in\mathcal{U}|{\hat{y}_{ul}}>{\hat{y}_{ui}},\forall i\in\mathcal{I}/l\}$
@@
Given the alignment of $\mathbf{r}$ and $\mathbf{q}_{1}$, and the pow-law distribution of the popularity, for any item $i\in\mathcal{I}/l$, the l.h.s of Eq. (20) can be bounded by:
(21) ${\sigma_{1}}{p_{1u}}({q_{1l}}-{q_{1i}})={\sigma_{1}}{p_{1u}}\frac{{{r_{l}}-{r_{i}}}}{{||{\mathbf{r}}||}}\geq{\sigma_{1}}{p_{1u}}\frac{{1-{2^{-\alpha}}}}{{\sqrt{\zeta(2\alpha)}}}$
@@
(22) $\sum\limits_{k=2}^{L}{{\sigma_{k}}{p_{ku}}({q_{ki}}-{q_{kl}})}\leq\sum\limits_{k=2}^{L}{\sqrt{2}{\sigma_{k}}}$
@@
(24) ${\sigma_{1}}{p_{1u}}\frac{{1-{2^{-\alpha}}}}{{\sqrt{\zeta(2\alpha)}}}>\sum\limits_{k=2}^{L}{\sqrt{2}{\sigma_{k}}}$
@@
(25) $\eta$ $=\frac{1}{n}|\{u\in U|{y_{ul}}>{y_{ui}},\forall i\in\mathcal{I}/l\}|$
@@
where $\phi(a)=\sum_{u\in\mathcal{U}}{\mathbf{I}[p_{1u}>a]}$ is an inverse cumulative function calculating the number of elements $p_{1u}$ in the left principal singular vector $\mathbf{p}_{1}$ exceeding a given value $a$, and the function $\mathbf{I}[.]$ signifies an indicator function.
A.3. Theoretical Analysis on Dimension Collapse via Gradient Optimization
Here we begin by invoking the gradient dynamic theorem from (Chou et al., 2023; Saxe et al., 2019; Arora et al., 2019) to elucidate the dimension collapse phenomenon and then develop a theorem to illustrate how the singular values impact the popularity bias in recommendations.
Theorem 1 (Trajectory of singular values (Eq. (6) in (Saxe et al., 2019))).
When training an MF model via gradient flow (gradient descent with infinitesimally small learning rate), i.e., $\frac{d}{{dt}}{\mathbf{U}}(t)=-\frac{{dL}}{{d{\mathbf{U}}}}$, $\frac{d}{{dt}}{\mathbf{V}}(t)=-\frac{{dL}}{{d{\mathbf{V}}}}$, the trajectory of singular values during the learning process obeys:
(26) $\sigma_{k}(t)=\frac{s_{k}e^{2s_{k}t}}{e^{2s_{k}t}-1+s_{k}/\sigma_{k}(0)}$